# Streaming Federated Learning with Markovian Data

### v1

@@ Appendix D Minibatch SGD
For the problem class $\mathcal{F}_{1}(L,\sigma,\tau)$, with constant step size $\gamma\leq\frac{1}{L}$, the iterates of Minibatch SGD satisfy:
$\Delta_{t+1}\leq\Delta_{t}-\frac{\gamma}{2}\mathbb{E}\left[\left\lVert\nabla F(w_{t})\right\rVert^{2}\right]+\frac{\gamma\tau\sigma^{2}}{MK}+\frac{\gamma\sigma^{2}}{2}\left\lVert\mathcal{Q}P^{Kt}-\pi\right\rVert_{TV}.$ (75)
@@
Note that the term $-\gamma\langle\nabla F\left(w_{t}\right),\bar{g}_{t}\rangle$ can be expanded as
@@
Plugging (77) back into (76), subtracting $F^{*}$ from both sides, and taking expectation we then have
$\Delta_{t+1}\leq\Delta_{t}-\frac{\gamma}{2}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right]+\frac{\gamma}{2}\mathbb{E}\left[\left\lVert\bar{g}_{t}-\nabla F\left(w_{t}\right)\right\rVert^{2}\right]-\left(\frac{1}{2\gamma}-\frac{L}{2}\right)s_{t}.$ (79)
Then with the condition $\gamma\leq\frac{1}{L}$, the last term in (79) can be ignored. The lemma follows by replacing the second term in (79) using Lemma C.1. ∎
@@
$\begin{split}\mathbb{E}\left[\left\lVert\nabla F(\hat{w}_{T})\right\rVert^{2}\right]\leq\mathcal{O}\left(\frac{\Delta_{0}}{\gamma T}+\frac{\tau\sigma^{2}}{MK}+\frac{\sigma^{2}}{T}\sum_{t=0}^{T-1}\left\lVert\mathcal{Q}P^{Kt}-\pi\right\rVert_{TV}\right).\end{split}$ (80)
@@
The proof of Theorem D.2 follows directly from Lemma D.1 by rearranging the terms and taking the average from $t=0$ to $T-1$
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right]\leq\frac{2\Delta_{0}}{\gamma T}+\frac{2\tau\sigma^{2}}{MK}+\frac{\sigma^{2}}{T}\sum_{t=0}^{T-1}\left\lVert\mathcal{Q}P^{Kt}-\pi\right\rVert_{TV}.$ (81)
@@
Corollary 4.6 can be derived directly from Theorem D.2.
@@
  1. (i)
    The system-level Markov chain $X$ is stationary $\left(Q=\pi\right)$,
    $K\geq\mathcal{O}\left(\frac{\tau\sigma^{2}}{M\epsilon^{2}}\right),\quad T\geq\mathcal{O}\left(\frac{L\Delta_{0}}{\epsilon^{2}}\right).$ (82)
  2. (ii)
    The system-level Markov chain $X$ is non-stationary $\left(Q\neq\pi\right)$, with $\epsilon\leq\mathcal{O}\left(\sqrt{\frac{\sigma^{2}}{M}}\right)$,
    $K\geq\mathcal{O}\left(\frac{\tau\sigma^{2}}{M\epsilon^{2}}\right),\quad T\geq\mathcal{O}\left(\frac{L\Delta_{0}\vee\sigma^{2}}{\epsilon^{2}}\right).$ (83)
@@
In the case where the system-level Markov chain is stationary, since $\pi P=\pi$, the term $\sum_{t=0}^{T-1}\left\lVert\mathcal{Q}P^{Kt}-\pi\right\rVert_{TV}$ in Theorem D.2 disappears. Choosing
$\begin{split}K&\geq\frac{4\tau\sigma^{2}}{M\epsilon^{2}},\\ T&\geq\frac{4\Delta_{0}}{\gamma\epsilon^{2}}\geq\frac{4L\Delta_{0}}{\epsilon^{2}},\end{split}$
guarantees that $\mathbb{E}\left[\left\lVert\nabla F\left(\hat{w}_{T}\right)\right\rVert^{2}\right]\leq\epsilon^{2}$.
In the case where the system-level Markov chain is non-stationary, by applying Lemma C.2 to Theorem D.2, we have
$\mathbb{E}\left[\left\lVert\nabla F(\hat{w}_{T})\right\rVert^{2}\right]\leq\frac{2\Delta_{0}}{\gamma T}+\frac{2\tau\sigma^{2}}{MK}+\frac{2\tau\sigma^{2}}{TK\ln 4}+\frac{\sigma^{2}}{T}.$
Choosing
$\begin{split}K&\geq\frac{8\tau\sigma^{2}}{M\epsilon^{2}},\\ T&\geq\frac{8\left(L\Delta_{0}\vee\sigma^{2}\right)}{\epsilon^{2}},\end{split}$
with
$\epsilon\leq\sqrt{\frac{8\sigma^{2}}{M}}$ (84)
to ensure $K\geq\tau$. This yield:
$\mathbb{E}\left[\left\lVert\nabla F(\hat{w}_{T})\right\rVert^{2}\right]\leq\frac{3}{4}\epsilon^{2}+\frac{M\epsilon^{2}}{64\ln 4\left(L\Delta_{0}\vee\sigma^{2}\right)}\epsilon^{2}.$
Hence, in the high-precision regime with
$\epsilon\leq\sqrt{\frac{8\sigma^{2}}{M}}\leq\sqrt{\frac{16\ln 4\left(L\Delta_{0}\vee\sigma^{2}\right)}{M}},$
we conclude that $\mathbb{E}\left[\left\lVert\nabla F(\hat{w}_{T})\right\rVert^{2}\right]\leq\epsilon^{2}$. ∎
Appendix E Local SGD
@@
$\xi_{t}=\frac{1}{MK}\sum_{m,k}\mathbb{E}\left[\left\lVert w_{t}^{(m,k)}-w_{t}\right\rVert^{2}\right],$
@@
Lemma E.1.
If the constant step size satisfies
@@
$\Delta_{t+1}\leq\Delta_{t}-\frac{45\tilde{\eta}}{98}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right]+\frac{2\tilde{\eta}\tau\sigma^{2}}{MK}+\frac{40}{98}\frac{L\tilde{\eta}^{2}\left(\theta^{2}+\sigma^{2}\right)}{\delta^{2}}+\tilde{\eta}\sigma^{2}\left\lVert\mathcal{Q}P^{Kt}-\pi\right\rVert_{TV}.$ (86)
@@
$\begin{split}\Delta_{t+1}&\leq\Delta_{t}-\frac{\tilde{\eta}}{2}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right]+\frac{\tilde{\eta}}{2}\mathbb{E}\left[\left\lVert g_{t}-\nabla F\left(w_{t}\right)\right\rVert^{2}\right]-\left(\frac{1}{2\tilde{\eta}}-\frac{L}{2}\right)s_{t}\\ &\leq\Delta_{t}-\frac{\tilde{\eta}}{2}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right]+\tilde{\eta}\mathbb{E}\left[\left\lVert\bar{g}_{t}-\nabla F\left(w_{t}\right)\right\rVert^{2}\right]+\tilde{\eta}\mathbb{E}\left[\left\lVert g_{t}-\bar{g}_{t}\right\rVert^{2}\right].\end{split}$ (87)
@@
Using the sample-wise smoothness Assumption 4.2 yields
$\begin{split}\mathbb{E}\left[\left\lVert g_{t}-\bar{g}_{t}\right\rVert^{2}\right]&=\mathbb{E}\left[\left\lVert\frac{1}{MK}\sum_{m,k}\nabla f_{m}\left(w_{t}^{(m,k)};x_{t}^{(m,k)}\right)-\nabla f_{m}\left(w_{t};x_{t}^{(m,k)}\right)\right\rVert^{2}\right]\\ &\leq\frac{L^{2}}{MK}\sum_{m,k}\mathbb{E}\left[\left\lVert w_{t}^{(m,k)}-w_{t}\right\rVert^{2}\right]\\ &=L^{2}\xi_{t}.\end{split}$ (88)
where the triangle inequality is used in the last step.
@@
$\begin{split}\xi_{t}=&\frac{1}{MK}\sum_{m,k}\mathbb{E}\left[\left\lVert w_{t}^{(m,k)}-w_{t}\right\rVert^{2}\right]\\ =&\frac{1}{MK}\sum_{m,k}\mathbb{E}\left[\left\lVert\sum_{j=1}^{k}\eta\nabla f_{m}\left(w_{t}^{(m,j)};x_{t}^{(m,j)}\right)\right\rVert^{2}\right]\\ \leq&\frac{1}{MK}\sum_{m,k}\eta^{2}k\sum_{j=1}^{k}\mathbb{E}\left[\left\lVert\nabla f_{m}\left(w_{t}^{(m,j)};x_{t}^{(m,j)}\right)\right\rVert^{2}\right]\\ \leq&\frac{\tilde{\eta}^{2}}{MK^{2}}\sum_{m,k}\sum_{j=1}^{k}\mathbb{E}\left[\left\lVert\nabla f_{m}\left(w_{t}^{(m,j)};x_{t}^{(m,j)}\right)\right\rVert^{2}\right]\\ \leq&\frac{\tilde{\eta}^{2}}{MK^{2}}\sum_{m,k}\sum_{j=1}^{K}\mathbb{E}\left[\left\lVert\nabla f_{m}\left(w_{t}^{(m,j)};x_{t}^{(m,j)}\right)\right\rVert^{2}\right]\\ \leq&\frac{\tilde{\eta}^{2}}{MK}\sum_{m,j}\mathbb{E}\left[\left\lVert\nabla f_{m}\left(w_{t}^{(m,j)};x_{t}^{(m,j)}\right)\right\rVert^{2}\right]\\ \leq&\frac{2\tilde{\eta}^{2}}{MK}\sum_{m,j}\mathbb{E}\left[\left\lVert\nabla f_{m}\left(w_{t}^{(m,j)};x_{t}^{(m,j)}\right)-\nabla f_{m}\left(w_{t};x_{t}^{(m,j)}\right)\right\rVert^{2}\right]\\ &+\frac{2\tilde{\eta}^{2}}{MK}\sum_{m,j}\mathbb{E}\left[\left\lVert\nabla f_{m}\left(w_{t};x_{t}^{(m,j)}\right)\right\rVert^{2}\right]\\ \overset{(i)}{\leq}&\quad 2L^{2}\eta^{2}\xi_{t}+\frac{4\tilde{\eta}^{2}}{MK}\sum_{m,j}\mathbb{E}\left[\left\lVert\nabla f_{m}\left(w_{t};x_{t}^{(m,j)}\right)-\nabla F_{m}(w_{t})\right\rVert^{2}\right]\\ &+\frac{4\tilde{\eta}^{2}}{M}\sum_{m}\mathbb{E}\left[\left\lVert\nabla F_{m}\left(w_{t}\right)\right\rVert^{2}\right]\\ \overset{(ii)}{\leq}&2L^{2}\tilde{\eta}^{2}\xi_{t}+4\tilde{\eta}^{2}\left(\theta^{2}+\sigma^{2}\right)+4\tilde{\eta}^{2}\delta^{2}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right].\end{split}$ (89)
@@
In $(i)$ we use the sample-wise smoothness (Assumption 4.2) and the triangle inequality, and in $(ii)$ we use the uniform bound on the noise of the gradient estimator (Assumption 4.3) and the heterogeneity assumption (Assumption 4.4).
Therefore we have:
@@
$\begin{split}\xi_{t}\leq\frac{4\tilde{\eta}^{2}\left(\theta^{2}+\sigma^{2}\right)}{1-2L^{2}\tilde{\eta}^{2}}+\frac{4\tilde{\eta}^{2}\delta^{2}}{1-2L^{2}\tilde{\eta}^{2}}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right].\end{split}$ (90)
@@
Plugging (88) and (90) back into (87), together with Lemma C.1 and the condition on the step size then yields
@@
$\begin{split}\Delta_{t+1}\leq&\Delta_{t}-\frac{45\tilde{\eta}}{98}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right]\\ &+\frac{2\tilde{\eta}\tau\sigma^{2}}{MK}+\frac{40}{98}L\tilde{\eta}^{2}\frac{\left(\theta^{2}+\sigma^{2}\right)}{\delta^{2}}+\tilde{\eta}\sigma^{2}\left\lVert\mathcal{Q}P^{Kt}-\pi\right\rVert_{TV},\end{split}$ (91)
which concludes the proof of the lemma. ∎
@@
Theorem E.2 (Theorem 4.7 in the main paper).
@@
$\mathbb{E}\left[\left\lVert\nabla F\left(\hat{w}_{T}\right)\right\rVert^{2}\right]\leq\mathcal{O}\left(\frac{\Delta_{0}}{\eta KT}+\frac{\tau\sigma^{2}}{MK}+\frac{L\eta K\left(\theta^{2}+\sigma^{2}\right)}{\delta^{2}}+\frac{\sigma^{2}}{T}\sum_{t=0}^{T-1}\left\lVert\mathcal{Q}P^{Kt}-\pi\right\rVert_{TV}\right).$ (93)
@@
By rearranging the terms in Lemma E.1 and taking the average from $t=0$ to $T-1$:
$\begin{split}\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right]&\leq\frac{98}{45}\frac{\Delta_{0}}{\tilde{\eta}T}+\frac{196}{45}\frac{\tau\sigma^{2}}{MK}+\frac{40}{45}\frac{L\tilde{\eta}\left(\theta^{2}+\sigma^{2}\right)}{\delta^{2}}+\frac{98}{45}\frac{\sigma^{2}}{T}\sum_{t=0}^{T-1}\left\lVert\mathcal{Q}P^{Kt}-\pi\right\rVert_{TV}.\\ \end{split}$ (94)
@@
Corollary E.3 (Corollary 4.8 in the main paper).
@@
the required number of local steps and communication rounds for Local SGD to achieve $\mathbb{E}[\|\nabla F(\hat{w}_{T})\|^{2}]\leq\epsilon^{2}$ are given by
  1. (i)
    If the system-level Markov chain $X$ is stationary $\left(Q=\pi\right)$,
    $K$ $\geq\mathcal{O}\left(\frac{\tau\sigma^{2}}{M\epsilon^{2}}\right),$ (96)
    $T$ $\geq\mathcal{O}\left(\frac{L\Delta_{0}}{\epsilon^{2}}\left(\delta^{2}\vee\frac{\theta^{2}+\sigma^{2}}{\delta^{2}\epsilon^{2}}\right)\right).$ (97)
  2. (ii)
    If the system-level Markov chain $X$ is non-stationary $\left(Q\neq\pi\right)$, with $\epsilon\leq\mathcal{O}\left(\sqrt{\frac{\sigma^{2}}{M}}\right)$,
    $K$ $\geq\mathcal{O}\left(\frac{\tau\sigma^{2}}{M\epsilon^{2}}\right),$ (98)
    $T$ $\geq\mathcal{O}\left(\frac{L\Delta_{0}\vee\sigma^{2}}{\epsilon^{2}}\left(\delta^{2}\vee\frac{\theta^{2}+\sigma^{2}}{\delta^{2}\epsilon^{2}}\right)\right).$ (99)
@@
In the case where the system-level Markov chain is stationary, since $\pi P=\pi$, the term $\sum_{t=0}^{T-1}\left\lVert\mathcal{Q}P^{Kt}-\pi\right\rVert_{TV}$ disappears. Choosing
$\begin{split}K&\geq\frac{196}{15}\frac{\tau\sigma^{2}}{M\epsilon^{2}},\\ \eta&\leq\left(\frac{3}{8}\frac{\delta^{2}\epsilon^{2}}{KL\left(\theta^{2}+\sigma^{2}\right)}\wedge\frac{1}{10KL\delta^{2}}\right),\\ T&\geq\frac{98}{15}\frac{\Delta_{0}}{\eta K\epsilon^{2}}\geq\frac{98}{15}\frac{L\Delta_{0}}{\epsilon^{2}}\left(10\delta^{2}\vee\frac{8}{3}\frac{\theta^{2}+\sigma^{2}}{\delta^{2}\epsilon^{2}}\right),\end{split}$
and applying Theorem E.2, then yields $\mathbb{E}\left[\left\lVert\nabla F\left(\hat{w}_{T}\right)\right\rVert^{2}\right]\leq\epsilon^{2}$.
In the case where the system-level Markov chain is non-stationary, by applying Lemma C.2 to Theorem E.2, we have
$\mathbb{E}\left[\left\lVert\nabla F(\hat{w}_{T})\right\rVert^{2}\right]\leq\frac{98}{45}\frac{\Delta_{0}}{\eta KT}+\frac{196}{45}\frac{\tau\sigma^{2}}{MK}+\frac{40}{45}\frac{L\eta K\left(\theta^{2}+\sigma^{2}\right)}{\delta^{2}}+\frac{98}{45}\frac{\sigma^{2}}{T}+\frac{196}{45}\frac{\tau\sigma^{2}}{KT\ln 4}.$
Choosing
$\begin{split}K&\geq\frac{196}{9}\frac{\tau\sigma^{2}}{M\epsilon^{2}},\\ \eta&\leq\left(\frac{9}{40}\frac{\delta^{2}\epsilon^{2}}{KL\left(\theta^{2}+\sigma^{2}\right)}\wedge\frac{1}{10KL\delta^{2}}\right),\\ T&\geq\frac{98}{9}\frac{\Delta_{0}\vee\sigma^{2}}{\eta K\epsilon^{2}}\geq\frac{98}{9}\frac{L\Delta_{0}\vee\sigma^{2}}{\epsilon^{2}}\left(10\delta^{2}\vee\frac{40}{9}\frac{\theta^{2}+\sigma^{2}}{\delta^{2}\epsilon^{2}}\right),\end{split}$
with
$\epsilon\leq\sqrt{\frac{196}{9}\frac{\sigma^{2}}{M}}$ (100)
to ensure that $K\geq\tau$. It follows that
$\mathbb{E}\left[\left\lVert\nabla F(\hat{w}_{T})\right\rVert^{2}\right]\leq\frac{4}{5}\epsilon^{2}+\frac{9}{490\ln 4}\frac{M\epsilon^{4}}{\left(L\Delta_{0}\vee\sigma^{2}\right)\left(10\delta^{2}\vee\frac{40}{9}\frac{\theta^{2}+\sigma^{2}}{\delta^{2}\epsilon^{2}}\right)}.$
In the high-precision regime,
$\epsilon\leq\sqrt{\frac{98\ln 4}{9}\frac{\sigma^{2}}{M}}\leq\sqrt{\frac{98\ln 4}{9}\frac{\sigma^{2}}{M}\left(10\delta^{2}\vee\frac{40}{9}\frac{\theta^{2}+\sigma^{2}}{\delta^{2}\epsilon^{2}}\right)},$
it then follows that $\mathbb{E}\left[\left\lVert\nabla F(\hat{w}_{T})\right\rVert^{2}\right]\leq\epsilon^{2}$. ∎
Appendix F Local SGD with Momentum

### v2

@@ D.1 Minibatch SGD
For the problem class $\mathcal{F}_{1}(L,\sigma,\tau)$, with constant global step size $\gamma\leq\frac{1}{L}$, the iterates of Minibatch SGD satisfy:
$\Delta_{t+1}\leq\Delta_{t}-\frac{\gamma}{2}\mathbb{E}\!\left[\left\lVert\nabla F(w_{t})\right\rVert^{2}\right]+\frac{2\gamma C_{\infty}\sigma^{2}}{\nu_{ps}MK}$ (79)
@@
Note that the term $\gamma\langle\nabla F\left(w_{t}\right),\bar{g}_{t}\rangle$ can be expanded as
@@
Plugging (81) back into (80), subtracting $F^{*}$ from both sides, and taking the conditional expectation, we then have
$\mathbb{E}_{t}\!\left[F(w_{t+1}-F^{*}\right]$ $\leq\mathbb{E}_{t}\!\left[F(w_{t})-F^{*}\right]-\frac{\gamma}{2}\mathbb{E}_{t}\!\left[\left\lVert\nabla F(w_{t})\right\rVert^{2}\right]+\frac{\gamma}{2}\mathbb{E}_{t}\!\left[\left\lVert\bar{g}_{t}-\nabla F(w_{t})\right\rVert^{2}\right]$ (83)
$-\left(\frac{1}{2\gamma}-\frac{L}{2}\right)\mathbb{E}_{t}\!\left[\left\lVert w_{t+1}-w_{t}\right\rVert^{2}\right].$ (84)
Then, with the condition $\gamma\leq\frac{1}{L}$, the last term in (83) can be ignored. The lemma follows by replacing the second term in (83) using Lemma D.1, then taking the full expectation. ∎
@@
$\mathbb{E}\!\left[\left\lVert\nabla F(\hat{w}_{t})\right\rVert^{2}\right]\leq\mathcal{O}\left(\frac{\Delta_{0}}{\gamma T}+\frac{C_{\infty}\sigma^{2}}{MK}\right)$ (85)
@@
The proof of Theorem D.3 follows directly from Lemma D.2 by rearranging the terms, then taking the average from $t=0$ to T1T-1
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right]\leq\frac{2\Delta_{0}}{\gamma T}+\frac{4C_{\infty}\sigma^{2}}{\nu_{ps}MK}$ (86)
@@
Now we are ready to prove Corollary 4.6.
@@
$K=\mathcal{O}\left(\frac{C_{\infty}\sigma^{2}}{\nu_{ps}M\epsilon^{2}}\right),\quad T=\mathcal{O}\left(\frac{L\Delta_{0}}{\epsilon^{2}}\right)$ (87)
@@
By choosing
$\begin{split}K&\geq\frac{8C_{\infty}\sigma^{2}}{\nu_{ps}M\epsilon^{2}},\\ T&\geq\frac{4\Delta_{0}}{\gamma\epsilon^{2}}\geq\frac{4L\Delta_{0}}{\epsilon^{2}},\end{split}$
(86) guarantees that $\mathbb{E}\left[\left\lVert\nabla F\left(\hat{w}_{T}\right)\right\rVert^{2}\right]\leq\epsilon^{2}$. ∎
D.2 Local SGD
@@
$\xi_{t}=\frac{1}{MK}\sum_{m,k}\mathbb{E}_{t}\left[\left\lVert w_{t}^{(m,k)}-w_{t}\right\rVert^{2}\right],$
@@
Lemma D.5.
If the local step size satisfies
@@
$\Delta_{t+1}\leq\Delta_{t}-\frac{45\tilde{\eta}}{98}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right]+\frac{4\tilde{\eta}C_{\infty}\sigma^{2}}{\nu_{ps}MK}+\frac{40}{98}\frac{L\tilde{\eta}^{2}\left(\theta^{2}+\sigma^{2}\right)}{\delta^{2}}.$ (89)
@@
$\begin{split}\mathbb{E}_{t}\!\left[F(w_{t+1})-F^{*}\right]&\leq\mathbb{E}_{t}\!\left[F(w_{t})-F^{*}\right]-\frac{\tilde{\eta}}{2}\mathbb{E}_{t}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right]+\frac{\tilde{\eta}}{2}\mathbb{E}_{t}\left[\left\lVert g_{t}-\nabla F\left(w_{t}\right)\right\rVert^{2}\right]\\ &-\left(\frac{1}{2\tilde{\eta}}-\frac{L}{2}\right)\mathbb{E}_{t}\!\left[\left\lVert w_{t+1}-w_{t}\right\rVert^{2}\right]\\ &\leq\mathbb{E}_{t}\!\left[F(w_{t})-F^{*}\right]-\frac{\tilde{\eta}}{2}\mathbb{E}_{t}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right]+\tilde{\eta}\mathbb{E}_{t}\left[\left\lVert\bar{g}_{t}-\nabla F\left(w_{t}\right)\right\rVert^{2}\right]\\ &+\tilde{\eta}\mathbb{E}_{t}\left[\left\lVert g_{t}-\bar{g}_{t}\right\rVert^{2}\right].\end{split}$ (90)
@@
Using the sample-wise smoothness (Assumption 4.2) yields:
$\begin{split}\mathbb{E}_{t}\left[\left\lVert g_{t}-\bar{g}_{t}\right\rVert^{2}\right]&=\mathbb{E}_{t}\left[\left\lVert\frac{1}{MK}\sum_{m,k}\nabla f_{m}\left(w_{t}^{(m,k)};x_{t}^{(m,k)}\right)-\nabla f_{m}\left(w_{t};x_{t}^{(m,k)}\right)\right\rVert^{2}\right]\\ &\leq\frac{L^{2}}{MK}\sum_{m,k}\mathbb{E}_{t}\left[\left\lVert w_{t}^{(m,k)}-w_{t}\right\rVert^{2}\right]\\ &=L^{2}\xi_{t}.\end{split}$ (91)
where Jensen’s inequality is used in the last step.
@@
$\begin{split}\xi_{t}=&\frac{1}{MK}\sum_{m,k}\mathbb{E}_{t}\left[\left\lVert w_{t}^{(m,k)}-w_{t}\right\rVert^{2}\right]\\ =&\frac{1}{MK}\sum_{m,k}\mathbb{E}_{t}\left[\left\lVert\sum_{j=0}^{k-1}\eta\nabla f_{m}\left(w_{t}^{(m,j)};x_{t}^{(m,j)}\right)\right\rVert^{2}\right]\\ \leq&\frac{1}{MK}\sum_{m,k}\eta^{2}k\sum_{j=0}^{k-1}\mathbb{E}_{t}\left[\left\lVert\nabla f_{m}\left(w_{t}^{(m,j)};x_{t}^{(m,j)}\right)\right\rVert^{2}\right]\\ \leq&\frac{\tilde{\eta}^{2}}{MK^{2}}\sum_{m,k}\sum_{j=0}^{k-1}\mathbb{E}_{t}\left[\left\lVert\nabla f_{m}\left(w_{t}^{(m,j)};x_{t}^{(m,j)}\right)\right\rVert^{2}\right]\\ \leq&\frac{\tilde{\eta}^{2}}{MK^{2}}\sum_{m,k}\sum_{j=0}^{K-1}\mathbb{E}_{t}\left[\left\lVert\nabla f_{m}\left(w_{t}^{(m,j)};x_{t}^{(m,j)}\right)\right\rVert^{2}\right]\\ \leq&\frac{\tilde{\eta}^{2}}{MK}\sum_{m,j}\mathbb{E}_{t}\left[\left\lVert\nabla f_{m}\left(w_{t}^{(m,j)};x_{t}^{(m,j)}\right)\right\rVert^{2}\right]\\ \leq&\frac{2\tilde{\eta}^{2}}{MK}\sum_{m,j}\mathbb{E}_{t}\left[\left\lVert\nabla f_{m}\left(w_{t}^{(m,j)};x_{t}^{(m,j)}\right)-\nabla f_{m}\left(w_{t};x_{t}^{(m,j)}\right)\right\rVert^{2}\right]\\ &+\frac{2\tilde{\eta}^{2}}{MK}\sum_{m,j}\mathbb{E}_{t}\left[\left\lVert\nabla f_{m}\left(w_{t};x_{t}^{(m,j)}\right)\right\rVert^{2}\right]\\ \overset{(i)}{\leq}&\quad 2L^{2}\eta^{2}\xi_{t}+\frac{4\tilde{\eta}^{2}}{MK}\sum_{m,j}\mathbb{E}_{t}\left[\left\lVert\nabla f_{m}\left(w_{t};x_{t}^{(m,j)}\right)-\nabla F_{m}(w_{t})\right\rVert^{2}\right]\\ &+\frac{4\tilde{\eta}^{2}}{M}\sum_{m}\mathbb{E}_{t}\left[\left\lVert\nabla F_{m}\left(w_{t}\right)\right\rVert^{2}\right]\\ \overset{(ii)}{\leq}&2L^{2}\tilde{\eta}^{2}\xi_{t}+4\tilde{\eta}^{2}\sigma^{2}+\frac{4\tilde{\eta}^{2}}{M}\sum_{m}\mathbb{E}_{t}\!\left[\left\lVert\nabla F_{m}(w_{t})\right\rVert^{2}\right].\end{split}$ (92)
@@
In $(i)$ we use the sample-wise smoothness (Assumption 4.2) and the triangle inequality, and in $(ii)$ we use the uniform bound on the noise of the gradient estimator (Assumption 4.3).
Therefore, we have:
@@
$\begin{split}\xi_{t}\leq\frac{4\tilde{\eta}^{2}}{1-2L^{2}\tilde{\eta}^{2}}\left(\sigma^{2}+\frac{1}{M}\sum_{m}\mathbb{E}_{t}\!\left[\left\lVert\nabla F_{m}(w_{t})\right\rVert^{2}\right]\right).\end{split}$ (93)
@@
Plugging (91) and (93) back into (90), together with Lemma D.1 and the condition on the step size then yields
@@
$\mathbb{E}_{t}\!\left[F(w_{t+1})-F^{*}\right]$ $\leq\mathbb{E}_{t}\!\left[F(w_{t})-F^{*}\right]-\frac{\tilde{\eta}}{2}\mathbb{E}_{t}\!\left[\left\lVert\nabla F(w_{t})\right\rVert^{2}\right]+\frac{4\tilde{\eta}C_{\infty}\sigma^{2}}{\nu_{ps}MK}$ (94)
$+\frac{40\tilde{\eta}^{2}L}{98\delta^{2}}\left(\sigma^{2}+\frac{1}{M}\sum_{m}\mathbb{E}_{t}\!\left[\left\lVert\nabla F_{m}(w_{t})\right\rVert^{2}\right]\right).$ (95)
By taking the full expectation and using the heterogeneity assumption 4.4, we have that:
$\Delta_{t+1}$ $\leq\Delta_{t}-\frac{\tilde{\eta}}{2}\left(1-\frac{80}{98}\tilde{\eta}L\delta^{2}\right)\mathbb{E}\!\left[\left\lVert\nabla F(w_{t})\right\rVert^{2}\right]+\frac{4\tilde{\eta}C_{\infty}\sigma^{2}}{\nu_{ps}MK}$ (96)
$+\frac{40\tilde{\eta}^{2}L}{98}\left(\frac{\theta^{2}+\sigma^{2}}{\delta^{2}}\right)$ (97)
Using the condition on the local step size on the second term in the RHS of (96) proves the lemma. ∎
@@
Theorem D.6 (Theorem 4.7 in the main paper).
@@
$\mathbb{E}\left[\left\lVert\nabla F\left(\hat{w}_{T}\right)\right\rVert^{2}\right]\leq\mathcal{O}\left(\frac{\Delta_{0}}{\eta KT}+\frac{C_{\infty}\sigma^{2}}{\nu_{ps}MK}+\frac{L\eta K\left(\theta^{2}+\sigma^{2}\right)}{\delta^{2}}\right).$ (99)
@@
By rearranging the terms in Lemma D.5 and taking the average from $t=0$ to T1T-1:
$\begin{split}\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left[\left\lVert\nabla F\left(w_{t}\right)\right\rVert^{2}\right]&\leq\frac{98}{45}\frac{\Delta_{0}}{\tilde{\eta}T}+\frac{392}{45}\frac{C_{\infty}\sigma^{2}}{\nu_{ps}MK}+\frac{40}{45}\frac{L\tilde{\eta}\left(\theta^{2}+\sigma^{2}\right)}{\delta^{2}}.\end{split}$ (100)
@@
Corollary D.7 (Corollary 4.8 in the main paper).
@@
the required number of local steps and communication rounds for Local SGD to achieve $\mathbb{E}[\|\nabla F(\hat{w}_{T})\|^{2}]\leq\epsilon^{2}$ are given by:
$K=\mathcal{O}\left(\frac{\tau\sigma^{2}}{M\epsilon^{2}}\right),\quad T=\mathcal{O}\left(\frac{L\Delta_{0}}{\epsilon^{2}}\left(\delta^{2}\vee\frac{\theta^{2}+\sigma^{2}}{\delta^{2}\epsilon^{2}}\right)\right).$ (102)
@@
By choosing
$\begin{split}K&\geq\frac{392}{15}\frac{C_{\infty}\sigma^{2}}{\nu_{ps}M\epsilon^{2}},\\ \eta&\leq\left(\frac{3}{8}\frac{\delta^{2}\epsilon^{2}}{KL\left(\theta^{2}+\sigma^{2}\right)}\wedge\frac{1}{10KL\delta^{2}}\right),\\ T&\geq\frac{98}{15}\frac{\Delta_{0}}{\eta K\epsilon^{2}}\geq\frac{98}{15}\frac{L\Delta_{0}}{\epsilon^{2}}\left(10\delta^{2}\vee\frac{8}{3}\frac{\theta^{2}+\sigma^{2}}{\delta^{2}\epsilon^{2}}\right),\end{split}$
(100) guarantees that $\mathbb{E}\left[\left\lVert\nabla F\left(\hat{w}_{T}\right)\right\rVert^{2}\right]\leq\epsilon^{2}$.
D.3 Local SGD with Momentum